# How well do LLMs reason over tabular data, really?

### v1

@@ Abstract
Large Language Models (LLMs) excel in natural language tasks but less is known about their reasoning capabilities over tabular data. Prior analyses devise evaluation strategies that poorly reflect an LLM’s realistic performance on tabular queries. Moreover, we have a limited understanding of the robustness of LLMs towards realistic variations in tabular inputs. Therefore, we ask: can general-purpose LLMs reason over tabular data, really?, and focus on two questions 1) are tabular reasoning capabilities of general-purpose LLMs robust to real-world characteristics of tabular inputs, and 2) how can we realistically evaluate an LLM’s performance on analytical tabular queries?
Building on a recent tabular reasoning benchmark, we first surface shortcomings of its multiple-choice prompt evaluation strategy, as well as commonly used free-form text metrics such as SacreBleu and BERT-score. We show that an LLM-as-a-judge procedure yields more reliable performance insights and unveil a significant deficit in tabular reasoning performance of LLMs. We then extend the tabular inputs reflecting three common characteristics in practice: 1) missing values, 2) duplicate entities, and 3) structural variations. Experiments show that the tabular reasoning capabilities of general-purpose LLMs suffer from these variations, stressing the importance of improving their robustness for realistic tabular inputs.
@@ 1 Introduction
Surfacing the reasoning capabilities on tabular tasks is, however, not straightforward. Most studies adopt free-form text metrics, which hardly capture reliable reasoning accuracy due to different formatting of the ground-truth answers, particularly of analytical questions, versus long-form LLM-generated responses Ji et al. (2024); Xu et al. (2023). The alternative of forcing certain output formats Sui et al. (2024) yields a limited understanding of the open-form reasoning performance and is prone to parsing errors, while another alternative of including the ground-truth answer in a multiple-choice prompt Qiu et al. (2024) leaks ground-truth answers into the prompt compromising its reliability. Beyond realistic evaluation procedures, in order to use LLMs in a reliable manner for tabular reasoning tasks, it is important to understand how well they can handle the characteristics of tabular data inputs Cong et al. (2023); Singha et al. as encountered in practice.
@@
Second, through comprehensive analysis we answer the question: are LLMs robust to real-world characteristics of tabular inputs? Building on the TQA-Bench, we first improve the validity of the task queries and downscale tabular data inputs to gain more fine-grained insights. We then inspect the robustness of LLMs against characteristics of tabular inputs as commonly found in practice. Specifically, we formalize the following three characteristics: missing values, duplicate entities, and structural variations. We show that most LLMs are not as robust to, and insufficiently acknowledge the presence of, such quality issues or anticipated variations in the tabular inputs, highlighting the need for more robust models for tabular reasoning.
@@ 2 Related Work
The TQA-Bench Qiu et al. (2024) examines multi-table reasoning capabilities with LLMs over various query complexities. We complement the TQA-Bench by using it as a base for our evaluation, and integrating common properties in tabular data, such as missing values, to study the robustness of multi-table reasoning capabilities of LLMs. Similarly, Sui et al. (2024) considers structural understanding capabilities by evaluating the accuracy of LLMs in basic tasks such as row/column retrieval. The QATCH benchmark Papicchio et al. (2023) evaluates tabular representation learning models specialized for SQL-centric tasks and mainly focuses on SQL-based evaluations. While the QATCH benchmark considers enterprise-centric evaluation tasks and inputs, it does not surface robustness for real-world properties. Earlier work by Cong et al. (2023) formalizes and analyzes key properties of tabular data principled in the relational data model, such as column-order insignificance. In this work, we focus on assessing the robustness on similar properties in the reasoning capabilities through the LLMs’ generated responses. In this realm, Singha et al. , evaluate various LLMs on their tabular understanding capabilities under noisy tabular inputs and variations in formatting of tabular data in prompts. While their robustness assessments cover realistic characteristics such as permutations in column-order, we include more properties and assess more complex reasoning capabilities of LLMs.
@@ 3.2 TQA-Bench Revisions
While TQA-Bench evaluates tabular reasoning capabilities with relatively large and multiple tables resulting in context sizes from 8K to 128K tokens. Our preliminary experiments revealed significant challenges in reasoning capabilities already with smaller table sizes, motivating the downsizing of context sizes to 1K, 2K, 4K, 6K and 8K to obtain more granular insights. To do so, we employ the scaling method introduced by TQA-Bench to truncate and segment tables while preserving their structural and relational integrity Qiu et al. (2024).
@@ 4 Towards Reliable Evaluation of Tabular Reasoning Capabilities
Evaluating tabular reasoning, and generally free-form question answering, still is an open challenge Ji et al. (2024); Xu et al. (2023). While context and reasoning traces of LLM-generated answers are useful, they complicate evaluation when ground-truth answers are short and exact, as is the case in typical tabular tasks. Table 1 illustrates this issue for two example queries from the popular OTTQA dataset for table question answering Chen et al. (2020) along with their ground-truth answers. When the LLM-generated answers are evaluated against the short ground-truth answers by two free-form text metrics, SacreBleu Post (2018) and BERT-score Zhang* et al. (2020), their scores are inconclusive. For example, for the queries in Table 1, the SB score for a correct generated answer is higher for the first query (1.4) but lower than the correct generated answer for the second query (0.5) for which the incorrect generated answer is closer to the correct answer for the first query (1.0). The BERT-score (BS) reflects mainly textual similarity, and shows barely any differences for different numeric values included in the response: its value is 0.81 for an incorrect as well as a correct generated answer. In what follows, we study the reliability of these different metrics in detail.
@@
Using the LLM-generated answers and ground-truth answers, we inspect the reliability of two comonly used free-form text metrics: SacreBleu Post (2018) and BERT-score Zhang* et al. (2020). SacreBLEU is a standardized version of the BLEU score that measures n-gram overlap between generated and reference texts, while the BERT-score leverages BERT embeddings to compute similarity based on token-level semantic matching. Our analysis reveals that neither the BERT-score nor SacreBleu metric provide a reliable signal for evaluating the correctness of generated answers. To visualize the reliability of these scores, we used Kernel Density Estimation (KDE) to estimate their distributions. For BERT-score, due to tight clustering of values, the KDE can exceed 1, while the more dispersed SacreBLEU values result in lower KDE peaks. The distributions of scores for correct and incorrect LLM-generated answers, when compared with ground-truth answers, exhibit significant overlap making them indistinguishable (Figures 1(a) and 1(b)). The inseparability between these distributions illustrates the unsuitability of these metrics for evaluating the accuracy of long-form answers against concise ground-truth answers.
@@
Recently, LLMs have emerged as a useful evaluation metric for free-form text, termed as LLM-as-a-judge evaluation Zheng et al. (2023). This approach to be particularly well-suited for tabular reasoning evaluation, where generated answers are often long- and free-form text compared to, for example OTTQA, where answers are short text snippets or single (numeric) values. We propose using an LLM-as-a-judge for evaluating tabular reasoning through LLMs, as this allows us to keep the generation close to real-world usage, where users expect models to generate complete answers rather than forcing a single-valued answer (which does not always correspond to the ground-truth answer) or choose from predefined options as in TQA-Bench Qiu et al. (2024). Second, relying on forced answer formats –accommodating multiple-choice or string-matching– can be brittle, especially for smaller models that may produce slightly misformatted outputs or fail to follow constraint templates.
@@ 5.3 Real-world characteristics of tabular data
Due to incomplete data collection or errors during data entry, tables often contain missing values, leading to incomplete information Little and Rubin (2019). To reflect this in the data, we first identify the cells needed to generate the answers and randomly remove one of the relevant cells. We recalculate the ground-truth answer by, effectively, setting the missing value to 0. The LLM-as-a-judge evaluates two behaviors: the model’s ability to produce the correct answer despite the missing information (Accuracy), and its capacity to explicitly acknowledge the absence of relevant data (Acknowledgement). These criteria reflect how well the model navigates incomplete data –whether it can reason effectively with what’s available– and transparently communicate the limitations introduced by missing values.
@@ 5.4 Robustness to Realistic Variation in Tables
Table 4: Results of the Missing Value perturbation across three aggregation tasks (average, sum and entity lookups) with table size 4k. The results illustrate that LLMs generally struggle with missing values but often acknowledge the missing values in their responses.
@@ Appendix A Full Benchmark
Size Model Entity lookup Top Selection Average Count Difference Sum Correlation

### v2

@@ Abstract
Large Language Models (LLMs) excel in natural language tasks, but less is known about their reasoning capabilities over tabular data. Prior analyses devise evaluation strategies that poorly reflect an LLM’s realistic performance on tabular queries. Moreover, we have a limited understanding of the robustness of LLMs towards realistic variations in tabular inputs. Therefore, we ask: Can general-purpose LLMs reason over tabular data, really?, and focus on two questions 1) are tabular reasoning capabilities of general-purpose LLMs robust to real-world characteristics of tabular inputs, and 2) how can we realistically evaluate an LLM’s performance on analytical tabular queries?
Building on a recent tabular reasoning benchmark, we first surface shortcomings of its multiple-choice prompt evaluation strategy, as well as commonly used free-form text metrics such as SacreBleu and BERT-score. We show that an LLM-as-a-judge procedure yields more reliable performance insights and unveil a significant deficit in tabular reasoning performance of LLMs. We then extend the tabular inputs reflecting three common characteristics in practice: 1) missing values, 2) duplicate entities, and 3) structural variations. Experiments show that the tabular reasoning capabilities of general-purpose LLMs suffer from these variations, stressing the importance of improving their robustness for realistic tabular inputs. [^1] Code: github.com/trl-lab/tabular-robustness
@@ 1 Introduction
Surfacing the reasoning capabilities on tabular tasks is, however, not straightforward. Most studies adopt free-form text metrics, which hardly capture reliable reasoning accuracy due to different formatting of the ground-truth answers, particularly of analytical questions, versus long-form LLM-generated responses Ji et al. (2024); Xu et al. (2023). The alternative of forcing certain output formats Sui et al. (2024) yields a limited understanding of the open-form reasoning performance and is prone to parsing errors, while another alternative of including the ground-truth answer in a multiple-choice prompt Qiu et al. (2024) leaks ground-truth answers into the prompt compromising its reliability. Beyond realistic evaluation procedures, in order to use LLMs in a reliable manner for tabular reasoning tasks, it is important to understand how well they can handle the characteristics of tabular data inputs Cong et al. (2023); Singha et al. (2023) as encountered in practice.
@@
Second, through comprehensive analysis we answer the question: are LLMs robust to real-world characteristics of multi-table inputs? Building on the TQA-Bench, we first improve the validity of the task queries and downscale multi-table inputs to gain more fine-grained insights. We then inspect the robustness of LLMs against characteristics of tabular inputs as commonly found in practice. Specifically, we formalize the following three characteristics: missing values, duplicate entities, and structural variations. We show that most LLMs are not as robust to, and insufficiently acknowledge the presence of, such quality issues or anticipated variations in multi-table inputs, highlighting the need for more robust models for tabular reasoning.
@@ 2 Related Work
The TQA-Bench Qiu et al. (2024) examines multi-table reasoning capabilities with LLMs over various query complexities. We complement the TQA-Bench by using it as a base for our evaluation, and integrating common properties in tabular data, such as missing values, to study the robustness of multi-table reasoning capabilities of LLMs. Similarly, Sui et al. (2024) considers structural understanding capabilities by evaluating the accuracy of LLMs in basic tasks such as row/column retrieval. The QATCH benchmark Papicchio et al. (2023) evaluates tabular representation learning models specialized for SQL-centric tasks and mainly focuses on SQL-based evaluations. While the QATCH benchmark considers enterprise-centric evaluation tasks and inputs, it does not surface robustness for real-world properties. Earlier work by Cong et al. (2023) formalizes and analyzes key properties of tabular data principled in the relational data model, such as column-order insignificance. In this work, we focus on assessing the robustness on similar properties in the reasoning capabilities through the LLMs’ generated responses. In this realm, Singha et al. (2023), evaluate various LLMs on their tabular understanding capabilities under noisy tabular inputs and variations in formatting of tabular data in prompts. While their robustness assessments cover realistic characteristics such as permutations in column-order, we include more properties and assess more complex reasoning capabilities of LLMs.
@@ 3.2 TQA-Bench Revisions
While TQA-Bench evaluates multi-table reasoning capabilities with relatively large and multiple tables resulting in context sizes from 8K to 128K tokens. Our preliminary experiments revealed significant challenges in reasoning capabilities already with smaller table sizes, motivating the downsizing of context sizes to 1K, 2K, 4K, 6K and 8K to obtain more granular insights. To do so, we employ the scaling method introduced by TQA-Bench to truncate and segment tables while preserving their structural and relational integrity Qiu et al. (2024).
@@ 4 Towards Reliable Evaluation of Tabular Reasoning Capabilities
Evaluating multi-table reasoning, and generally free-form question answering, still is an open challenge Ji et al. (2024); Xu et al. (2023). While context and reasoning traces of LLM-generated answers are useful, they complicate evaluation when ground-truth answers are short and exact, as is the case in typical tabular tasks. Table 1 illustrates this issue for two example queries from the popular OTTQA dataset for table question answering Chen et al. (2020) along with their ground-truth answers. When the LLM-generated answers are evaluated against the short ground-truth answers by two free-form text metrics, SacreBleu Post (2018) and BERT-score Zhang* et al. (2020), their scores are inconclusive. For example, for the queries in Table 1, the SB score for a correct generated answer is higher for the first query (1.4) but lower than the correct generated answer for the second query (0.5) for which the incorrect generated answer is closer to the correct answer for the first query (1.0). The BERT-score (BS) reflects mainly textual similarity, and shows barely any differences for different numeric values included in the response: its value is 0.81 for an incorrect as well as a correct generated answer. In what follows, we study the reliability of these different metrics in detail.
@@
Using the LLM-generated answers and ground-truth answers, we inspect the reliability of two commonly used free-form text metrics: SacreBleu Post (2018) and BERT-score Zhang* et al. (2020). SacreBLEU is a standardized version of the BLEU score that measures n-gram overlap between generated and reference texts, while the BERT-score leverages BERT embeddings to compute similarity based on token-level semantic matching. Our analysis reveals that neither the BERT-score nor SacreBleu metric provide a reliable signal for evaluating the correctness of generated answers. To visualize the reliability of these scores, we used Kernel Density Estimation (KDE) to estimate their distributions. For BERT-score, due to tight clustering of values, the KDE can exceed 1, while the more dispersed SacreBLEU values result in lower KDE peaks. The distributions of scores for correct and incorrect LLM-generated answers, when compared with ground-truth answers, exhibit significant overlap making them indistinguishable (Figures 1(a) and 1(b)). The inseparability between these distributions illustrates the unsuitability of these metrics for evaluating the accuracy of long-form answers against concise ground-truth answers.
@@
Recently, LLMs have emerged as a useful evaluation metric for free-form text, termed as LLM-as-a-judge evaluation Zheng et al. (2023). This approach to be particularly well-suited for tabular reasoning evaluation, where generated answers are often long- and free-form text compared to, for example OTTQA, where answers are short text snippets or single (numeric) values. We propose using an LLM-as-a-judge for evaluating tabular reasoning through LLMs, as this allows us to keep the generation close to real-world usage, where users expect models to generate complete answers rather than forcing a single-valued answer (which does not always correspond to the ground-truth answer) or choose from predefined options as in TQA-Bench Qiu et al. (2024). Second, relying on forced answer formats –accommodating multiple-choice or string-matching– can be brittle, especially for smaller models that may produce slightly misformatted outputs or fail to follow constraint templates Liu et al. (2024).
@@ 5.3 Real-world characteristics of tabular data
Due to incomplete data collection or errors during data entry, tables often contain missing values, leading to incomplete information Little and Rubin (2019). This has been a longstanding issue for predictive ML, as missing values can distort analysis and lead to unreliable results Emmanuel et al. (2021). To reflect this in the data, we first identify the cells needed to generate the answers and randomly remove one of the relevant cells. We recalculate the ground-truth answer by, effectively, setting the missing value to 0. The LLM-as-a-judge evaluates two behaviors: the model’s ability to produce the correct answer despite the missing information (Accuracy), and its capacity to explicitly acknowledge the absence of relevant data (Acknowledgement). These criteria reflect how well the model navigates incomplete data—whether it can reason effectively with what’s available—and transparently communicate the limitations introduced by missing values.
@@ 5.4 Robustness to Realistic Variation in Tables
Table 4: Results of the Missing Value perturbation across three aggregation tasks (average, sum and entity lookups) with table size 4k.
@@ Appendix A Full Benchmark
Size Model Entity lookup Top Selection Average Count Subtraction Sum Correlation